OpenStudy (anonymous):

Explain and show how you would estimate zero point five times four point six two. Then justify why this is a good choice for arriving at the correct answer

OpenStudy (anonymous):

The method used for estimating zero point five times four point six two would be to divide four point six two by two. This would give the answer of two point three one. The reasoning behind this approach is that point five times anything is always half of that anything. Half of anything creates two equal parts. So dividing anything by two results in the same answer without needing to perform a long multiplication process. The problem as stated: (0.5) (4.62) = X Estimate: X=4.62/2 Work: 4/2=2, 6/2=3, 2/2=1 Answer: X=2.31

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

Welcome!!:)